6. Solve the differential equation by variation of parameters. y – 2y + y = fiz
Answer #1

Auxillary eght g - 2y + y = en (tx² m² _ amtlzo (may=0 I mali Ye (x) = + C₂x) en - Yg = a 4= e Whyudo = ex xen en entren xeera (1+x2yeza tan (2) a & Yp (2) e 2 lu (1+x2) + ne tan (2) 기 + y (2) 42) = ye(x) = (cit Cax) e + ne + ne tana - eu lu (1